OpenStudy (anonymous):

Find the approximate values for the derivative of f(x) at each of the x - values given x 0,1,2,3,4,5,6,7,8, f(x) 18, 13, 10, 7, 9, 9, 15, 19, 30

OpenStudy (dumbcow):

f'(0) = (13-18)/1-0 = -5 f'(1) = (10-13)/(2-1) = -3 ... i guess thats the best approximation you can do just given the points